Skip to main content

Daily News

Project Acronym: HSDPD | State: Completed | Expedition ID: 5053

▾▴ Toggle Texts
▾▴ Toggle Images
◱ Full Size Pics
⊻ Bottom of Page

... (?TBD)

Actions

Displaying results 51 to 79 out of 79
<< First< Previous1-5051-79Next >Last >>
  • (inserted: 07.07.2013 20:00)
    HSPDP: 2013-07-08

    Looks like drilling is finished at W Turkana. Communicated from ...


    Looks like drilling is finished at W Turkana. Communicated from Chris today. Yesterday we tried getting the HQ back down to 215m, but were stopped at about 200 meters (5 rods short). Couldn't latch on to the tool so tripped out the rods to find out that there were 6 meters of cuttings on TOP of the tool. Tried again today, and got back down to about 200m. Went to clear the tool, but again couldn't get to it - guessing the same thing happened as yesterday (DPI is currently tripping out HQ). We think that there is another aquifer around 200 m that is forcing cutting back up between the tool and the HQ, basically overpowering Mike's pump (which is rated to 1000 psi, but blows hoses at 300). We are going to go back into the casing, then try to get back to 200m again tomorrow in time for the loggers to log inside the HQ to at least that depth. Definitely no more core to get beyond 215m. Drilling Report: Yesterday we tried getting the HQ back down to 215m, but were stopped at about 200 meters (5 rods short). Couldn't latch on to the tool so tripped out the rods to find out that there were 6 meters of cuttings on TOP of the tool. Tried again today, and got back down to about 200m. We think that there is another aquifer around 200 m that is forcing cutting back up between the tool and the HQ, basically overpowering Mike's pump. We are going to go back into the casing, then try to get back to 200m again tomorrow in time for the loggers to log inside the HQ to at least that depth. Definitely no more core to get beyond 215m. So final depth of 215m and with any luck, we'll be able to log in the HQ from about 196m up (200 - 4m tool) and possibly open hole from about 135-150 m up where things seem to be a bit more stable if we can get the casing. So final depth of 215m and with any luck, we'll be able to log in the HQ from about 196m up (200 - 4m tool) and possibly open hole from about 135-150 m up where things seem to be a bit more stable if we can get the casing.
  • (inserted: 06.07.2013 20:00)
    HSPDP: 2013-07-07

    Drilling Report: Some good news! This morning we got the ...


    Drilling Report: Some good news! This morning we got the casing unstuck and advanced down to the level of the stuck casing (about 162m), felt it hit, then slide off and be redirected to the side, or so we believe. There is a possibility it is simply driving the stuck casing down, but it seemed to be moving much too easy for that to be the case. At about 4pm we finally got bogged down by reaming with the casing at about 183 or 186 m. The HQ rods are being tripped in (should be done) and we will start coring with long shifts again tomorrow at 7am and go until water is an issue, which should be at least 16 hours if not more based on prior days. Assuming no more major complications (knock on wood), we don't see any reason why with low torque we can't get some 30m a day with extended shifts. Ideally, we will at least get back to the 212 mark tomorrow, then 60 more meters on the 8th and 9th, putting us at about 275m depth. Craig and I did some calculations and we though 280-300m would possibly be sufficient for the full lake phase (and at least the bottom of the Olduvai), so we're crossing our fingers. We were knocked down, but back to fighting.
  • (inserted: 05.07.2013 20:00)
    HSPDP: 2013-07-06

    Drilling Report: This morning we got the casing unstuck and ...


    Drilling Report: This morning we got the casing unstuck and advanced down to the level of the stuck casing (about 162m), felt it hit, then slide off and be redirected to the side, or so we believe. There is a possibility it is simply driving the stuck casing down, but it seemed to be moving much to easy for that to be the case. At about 4pm we finally got bogged down by reaming with the casing at about 183 or 186 m (Mike and Rich have different casing counts I think). The HQ rods are being tripped in (should be done) and we will start coring with long shifts again tomorrow at 7am and go until water is an issue, which should be at least 16 hours if not more based on prior days. Assuming no more major complications (knock on wood), we don't see any reason why with low torque we can't get some 30m a day with extended shifts. Ideally, we will at least get back to the 212 mark tomorrow, then 60 more meters on the 8th and 9th, putting us at about 275m depth. Craig and I did some calculations and we though 280-300m would possibly be sufficient for the full lake phase (and at least the bottom of the Olduvai), so we're crossing our fingers. We were knocked down, but back to fighting.
  • (inserted: 04.07.2013 20:00)
    HSPDP: 2013-07-05

    The last few days have been very slow and a ...


    The last few days have been very slow and a bit frustrating for all. We are encountering a combination of problems that are impeding progress. First and most significantly, the drillers are experiencing very high torque on the turning rods, presumably the result of friction with excessive mud building up in the borehole. Attempts to clear this with a combination of bentonite drilling mud and polymer have been only marginally successful. On the 3rd we advanced about 12 meters but eventually it became impossible to go deeper.
  • (inserted: 30.06.2013 20:00)
    HSPDP: 2013-07-01

    The photo is of our nurse Gladys Tuitoek, who (we ...


    The photo is of our nurse Gladys Tuitoek, who (we are happy to report) has nothing to do on the medical front, so we have put her to work labeling cores instead-maybe she'll switch to being a geologist!
  • (inserted: 29.06.2013 20:00)
    HSPDP: 2013-06-30

    Things are moving ahead at the drill site. I'm back ...


    Things are moving ahead at the drill site. I'm back in the US now and Tom Johnson and Dan Olago left Kaitio on the 28th. I heard from Chris a short while ago. They were wrapping up for the day (9PM) with about 27m of core collected in 10 hours after a few days of casing problems and little drilling progress. Currently we are at 187m, so slightly over half the target depth. DPI plans to thicken the mud tomorrow morning-they have had to go to using polymer so geochem analysis of the lower part of the core will have to take this into consideration.
  • (inserted: 27.06.2013 20:00)
    HSPDP: 2013-06-28

    Communications have been quiet the last few days due to ...


    Communications have been quiet the last few days due to the difficulty of communicating from the drill site in Northern Kenya. After a day and a half of slow progress, tripping pipe and casing we are back in business drilling! Down to 155m. Our 3-D film team from Earth Images Foundation was with us the last few days and left yesterday. We are very excited about the film they will produce of the project.
  • (inserted: 24.06.2013 20:00)
    HSPDP: 2013-06-25

    Drilling Report: Day 3 of drilling with still another 12 ...


    Drilling Report: Day 3 of drilling with still another 12 hour day but ended at 104m with excellent recovery! The 45 meters drilled produced a very long sequence of lacustrine mudstone.
  • (inserted: 23.06.2013 20:00)
    HSPDP: 2013-06-24

    Yesterday was the start of drilling. It was a slow ...


    Yesterday was the start of drilling. It was a slow start with a 12 hour drilling day ending at 22m. Today was another 12 hour day but drilling speed increased ending the day with 37m drilled and close to a 100% core recovery. Our goal is to ramp up the drilling time to 15 hours then 24 hours as soon as possible. Drilling Report: Day 2 of drilling. We had another 12 hour shift but got about 37m of core with almost 100% recovery. We intend to ramp it up to 15 hours of drilling then to 24 hours.
  • (inserted: 22.06.2013 20:00)
    HSPDP: 2013-06-23

    Drilling Report: It was a slow set up. We had ...


    Drilling Report: It was a slow set up. We had a 12 hour drilling day with 3m of over burden then 15m of core recovery ending at 22m depth. Core recovery improved after advancing to 18m casing.
  • (inserted: 17.06.2013 20:00)
    HSPDP: 2013-06-18

    With the exception of a few personnel who are in ...


    With the exception of a few personnel who are in Kenya for both phases, the Tugen Hills crew has left Kenya for the most part. The Turkana crew is all set and flying up to Lodwar on the 19th, then driving to the drill site on the 20th. The drill rig and equipment will arrive about the same time and our mobile field camp is already set-up and ready to go. As cell communications are non-existent at the drill/camp site, we'll be relying on (expensive) satellite communications for Turkana. As such, updates may be a bit more sporadic as we send them out with people coming and going from camp.
  • (inserted: 14.06.2013 20:00)
    HSPDP: 2013-06-15

    The Baringo Tugen Hills drilling site team returned from the ...


    The Baringo Tugen Hills drilling site team returned from the field to Nairobi yesterday. Everyone was in a great mood with the successful drilling campaign behind us. The West Turkana team is now assembling. We will pack the freight truck on Monday and then we fly up to Lodwar on Wednesday in time to meet the truck and other members of the team (Craig Feibel and Kat Beck) who are already in north Kenya.
  • (inserted: 12.06.2013 20:00)
    HSPDP: 2013-06-13

    The Tugen Hills drilling finished up on June 11th. The ...


    The Tugen Hills drilling finished up on June 11th. The downhole logging continued and has been completed. We reached a depth of 228 meters, just shy of our original 250m target depth, but we may have reached further back in time than originally anticipated. If so, this will provide even more overlap for a great comparison to the northern Awash core. The downhole logging is also complete and the science crew is back in Nairobi. In the next few days we will be swapping out some project personnel and heading up to West Turkana on the 19th to begin the second phase of drilling. A festive evening capped off the end of the Tugen Hills drilling, which included some traditional dancers. So far, so good. Stay tuned for more updates from Nairobi and West Turkana!
  • (inserted: 11.06.2013 20:00)
    HSPDP: 2013-06-12

    The ICDP downhole logging crew arrived at the Tugen-Hill drill ...


    The ICDP downhole logging crew arrived at the Tugen-Hill drill site, started to rig up the logging unit and prepared the slimhole sonds for the first logging runs.
  • (inserted: 10.06.2013 20:00)
    HSPDP: 2013-06-11

    The deeper we go, the more challenging the drilling gets ...


    The deeper we go, the more challenging the drilling gets - formation swelling, sticking rods, heavy rain and lightning. Despite the resistance of the Chemeron Formation, we are down to about 228 meters. During a break in drilling, His Excellency, the Honorable Benjamin C. Cheboi, Governor of Baringo County, visited the project at the drill site and in the laboratory at Soi Lodge.
  • (inserted: 08.06.2013 20:00)
    HSPDP: 2013-06-09

    Drilling Report: Arrived on site and began tripping rods back ...


    Drilling Report: Arrived on site and began tripping rods back to bottom. We hit a bridge around 150m and began to ream the rest of the rods down to bottom. Beau and I cleaned up the new inner tubes, brushing out all the rust build up along with cleaning up the threads so that the new liners would slide in easier. We began coring at 205.25m and through the first run it was apparent that the rig was bogging down due to the formation swelling. Torque is very high and penetration is very difficult. We tried making a thick batch of mud to try and lift all the cuttings out; that helped a little. To save on Bentonite consumption we changed the mud mix to a Polly/water/Bentonite mix and that seemed to help a lot. Blew two hydraulic hoses but were able to get them fixed. After drilling a few runs we ended the shift at 210.99m. Rods were sticking all day and penetration rate very slow, we needed to trip the rod back up to the problem area between 150m and 180m and ream the hole to try and condition it. Had some heavy rain and lightning come in late in the shift, we shut down at 11pm and will resume operations in the morning.
  • (inserted: 07.06.2013 20:00)
    HSPDP: 2013-06-08

    Drilling Report: Started the day at 204.85 and had to ...


    Drilling Report: Started the day at 204.85 and had to core 0.4m to allow access for HW casing. Tried to rotate and lift casing to drive down further. Casing was stuck and not able to advance casing without compromising stability of borehole. Tripped out HQ rods to inspect bit. With only new liners on hand they are slightly larger than the previous liners and do not fit into the inner tube. Standby 12:00pm to 6pm waiting for Beau to arrive with new inner tubes from Nairobi to accommodate the new liners. Beau and inner tubes arrived late after a long day of travel and we will start first thing in the morning and resume normal drilling operations.
  • (inserted: 06.06.2013 20:00)
    HSPDP: 2013-06-07

    Over the last few days we had a variety of ...


    Over the last few days we had a variety of small glitches. Thursday evening we got our first significant rain at the drill site and saw the effect on our road in and out. There is a big hill coming down from the drill site and when the truck reached it, it started sliding even in 4WD-low. We parked it and called for the other vehicle to come from the lodge to get us but when the second car parked at the base of the hill we could not get it started again! So now we had two stuck vehicles and a 2km walk out to the main road where the hotel mini-van picked us up. A new load of core liner was delivered, but when we tried putting these into the coring tool inner tubes they wouldn't fit! The upshot is that Beau Marshall arrived in Nairobi late last night and went to the NOCK yard where 3 more inner tubes were in the DES containers and found that the liner does fit into those inner tubes. On the positive side, the Multisensor Core Logger is now up and running at the Soi Hotel and we are getting very nice data. The diatomites from deep-water lake phases show up very nicely. MSCL logging is moving ahead at about 10m an hour, so if nothing happens (fingers crossed against blackouts!) we will clear out the backlog of cores by tomorrow afternoon. Finally, we are currently at 204m, so over 80% of the way to our 250 target depth. Al Deino thinks we will have a significant temporal overlap in this record with the Afar record, which would be really great. Drilling Report: With one vehicle broken down in Nairobi and one broken down at the bottom of the hill we took a taxi to the bottom of the hill and then drove the Land Cruiser that was at the top of the hill to the site. Starting the day at 189.6m, both drill crews working together on the same shift we began coring. Half way through the second run the main hydraulic hose blew at 10:30am we had to shut down and go get a new hose made up. Drilling resumed at 4:30pm and cored down to 204.85m for a total of 15.25m. But with the torque climbing as we got deeper and the runs taking longer we stopped drilling at 11:30pm and will advance the casing in the morning.
  • (inserted: 04.06.2013 20:00)
    HSPDP: 2013-06-05

    Drilling Report: Started coring at 119.40m, with Mike in Nairobi ...


    Drilling Report: Started coring at 119.40m, with Mike in Nairobi getting his pins taken out of his leg today we are still on one extended drilling shift with both DPI crews working together while Mike is gone. We are still having issues with the tube getting stuck all day long with the sand/silt buildup in the rods. We have drained the return sump and turned it into and holding tank for clean fresh mud, this way we will not run out of mud during a drill run. We are no longer going to recirculate the mud coming out of the casing. (Made a trench to funnel it into the brush). We ended the shift at 11pm and finished coring at 156.0m with at total of 36.6m and 91% recovery.
  • (inserted: 03.06.2013 20:00)
    HSPDP: 2013-06-04

    110 meters of core has been recovered (ahead of schedule) ...


    110 meters of core has been recovered (ahead of schedule) with 98% recover over the last 60m! Varied lithology reflecting oscillating lake transgression/regression sequences with abundant tuffs. Density profiles of cores/depth in hole neatly reflect cycling of diatomites in response to eccentricity modulated precession. Currently collecting interesting facies indicating varied terrestrial and potentially lacustrine environmental conditions.
  • (inserted: 02.06.2013 20:00)
    HSPDP: 2013-06-03

    Drilling Report: After night shift advanced the casing we are ...


    Drilling Report: After night shift advanced the casing we are now getting about 90% returns. Day shift started coring at 46.65m and ending at 73.70 with a total of 27.05m cored (98% recovery). Day shift also had one stuck tube due to sticky clay/sand but managed to get it free after working it for a while. Start of Night shift cored from 73.70 and ending at 88.95 with a total of 15.25m cored (98% recovery) but having to trip out the rods twice to get the tool out of the core barrel. We are getting a fine sandy/clay like wall cake on the inside of the rods. We are having the drill crews flush the rods before each run to try and wash them down. Total cored drilled for entire day was 42.30m with a 98% recovery.
  • (inserted: 31.05.2013 20:00)
    HSPDP: 2013-06-01

    First day of drilling. A single 12 hour daylight shift ...


    First day of drilling. A single 12 hour daylight shift with all crew involved. Successfully drilled 25 meters! We will now begin 24 hour drilling with two 12-hour shifts. Photos above are of the Tugen Hills science team and the first Tugen Hills core (3 meters) on deck. Drilling Report: Breakfast at 6:30 then all members drive to site for a single 12 hour shift with both day and night shifts working together to get everyone up to speed and trained. 7:45 science team setting up their tent along with other various equipment. 8:30 sight orientation with all members of science and drill crew to go over safety topics and procedures along with general drilling operations. Start coring at 9:15 and drilled to a final depth of 24.90m for the shift. At the end of the shift DPI tripped out to change to a different style bit and shoe to try and improve core recovery for the day shift in the morning. Great first day for everyone!!!
  • (inserted: 29.05.2013 20:00)
    HSPDP: 2013-05-30

    The shipping containers finally made it from the docks in ...


    The shipping containers finally made it from the docks in Mombasa to the NOCK (National Oil Corporation of Kenya) yard. Unfortunately, the truck was a bit too big for the NOCK entry gate! After removing some keypad posts and gate arms, the truck and crane finally made it in to be unloaded. Gear to be transported to the awaiting drill rig on May 31 with drilling planned for June 1. Drilling Report: Leave Norfolk Towers at 8am (Anders, Beau and Rich) and take taxi to NOCK yard. Wait on crane and containers. Truck with containers arrive shortly before 9am and waited on crane until 9:45ish. Entrance gate to NOCK yard extremely tight and difficult to navigate equipment thru. Crane managed to squeeze its way in, semi truck with containers ate up some time. Had to remove 2 electronic key pad authorization posts as well as electronic gate arm. Crane set up and containers were finally on the ground by 12noon. The next 1.5 hours we unloaded items from containers and into the land rover. Headed back to apartment. Due to the delays encountered this morning, we pushed the travel to Baringo back until tomorrow morning.DPI crews continue with minor rigging up tasks, ready to drill, standby and wait on science crews to arrive. Drilling still on schedule to start Saturday June 1.
  • (inserted: 28.05.2013 20:00)
    HSPDP: 2013-05-29

    Drilling Report: After breakfast at Baringo Mike picked up Beau ...


    Drilling Report: After breakfast at Baringo Mike picked up Beau and I, then headed to Nairobi so we could receive the containers that afternoon. On the ride there Beau was communicating with the shipping agent and truck driver that would be delivering them. After talking to the driver he informed Beau that he would not be able to unload the containers until 9am the following morning. We dropped Mike off at his house and then met up with the science team at Norfolk Towers and discussed some project issues. Beau and I will be at Nock Yard at 9am to unload supplies needed for the project and load them into our vehicles. DPI crew finished digging the circulation pits, lined and filled them with water also doing any other minor last minute rigging up. The DPI crew and site are ready to start drilling.
  • (inserted: 27.05.2013 20:00)
    HSPDP: 2013-05-28

    Drilling Report: Arrived on site with DPI already starting to ...


    Drilling Report: Arrived on site with DPI already starting to dig out the circulation pits. DPI trucks (two) traveled to Nairobi to get water storage tanks. Measured tools inside core barrels to see if any adjustments need to be made. Hoisted up core barrel with tool inside and lowered down inside the casing to bottom with mainline to see if any debris or fill was inside of casing, rods went all the way down to bottom, casing is clear. DPI crew set up lighting system for night drilling operations, fuel/flammables containment area and will finish digging circulation pits in the morning, line pits and fill with water. Aside from minor rigging up and finishing touches, we are pretty much ready to start drilling operations on June 1st 2013. Beau communicated with Judy (Import Tracking) at Kenfreight and received confirmation that the containers will be loaded on trucks today. Just after 8pm he received confirmation that trucks are leaving Mombasa in the morning and will arrive in Nairobi (NOCK Yard) sometime tomorrow afternoon. Beau and I will be traveling back to Nairobi at 7:30am tomorrow morning to receive the containers and unload items on Thursday. Beau flies out of Nairobi early Friday morning and Mike and I will travel back to Baringo Friday as well. Everything is on schedule!
  • (inserted: 26.05.2013 20:00)
    HSPDP: 2013-05-27

    The DPI drill rig is in place at the Tugen ...


    The DPI drill rig is in place at the Tugen Hills site. The trucks had no problem making it up the newly made/graded road to the site. The capping Kapthurin gravels have been drilled and cased so we are all set to drill the underlying Pliocene lacustrine sediments once the rest of the gear clears customs. Water pump set up in river about 50 meters below the site works great and should allow for 24/7 drilling no problem. Drilling Report: Set up rig. Connect 6 inch casing to ODEX hammer. Drill 0-5.25m (Kapthurin Gravel cover) to formation change. Remove hammer and grout casing in place. Connect water system from river using 220m of 2 inch pipe. Test pump. Note: good volume of drilling water from river. Start digging circulation pits. To continue digging circulation pits in the morning.
  • (inserted: 25.05.2013 20:00)
    HSPDP: 2013-05-26

    Mobilization. Drive from Marigat town to site. Marked borehole position ...


    Mobilization. Drive from Marigat town to site. Marked borehole position with John and Beau. Moved rig into position. Offloaded all trucks and set up camp. Will connect water system from river and dig circulation pits in the morning.
  • (inserted: 24.05.2013 20:00)
    HSPDP: 2013-05-25

    Met Mike Scarpelini with DPI after breakfast and started ...


    Met Mike Scarpelini with DPI after breakfast and started driving to Baringo. Stopped in Nakuru for a short while, continued journey to Baringo and arrived to turn off for the drill site. Drove to the base of the hill, then up the hill to test it. Looks to be a decent improvement. Head to Tamarind Lodge for the night. Joseph with Beacon Movers is difficult to get a hold of, cannot get confirmation on container delivery. Still waiting on containers. DPI trucks stopped for the night in Marigat, will leave early in the morning and meet us on the side of the road at the turn off to the drill site.
  • (inserted: 23.05.2013 20:00)
    HSPDP: 2013-05-24

    Beau called Joseph with Beacon Movers to get update on ...


    Beau called Joseph with Beacon Movers to get update on when containers would arrive in Nairobi. It appears they still have not been released in Mombasa and he said maybe Monday or Tuesday. Went and got local cellphone set up. Returned to apartment, worked on receipts, paperwork and other admin tasks. John Kingston went up to Baringo to check out road work and pinpoint drill hole. Andy Cohen and science team begin travel from USA to Nairobi. Plan to unload containers today did not happen, we will continue with the plan to mobilize to Baringo tomorrow with DPI. DPI loaded equipment onto trucks in their yard in preparation for mobilization tomorrow.
Displaying results 51 to 79 out of 79
<< First< Previous1-5051-79Next >Last >>

To top